# Canonical decompositions of 3-connected graphs

### v2

@@ 2.7. Tutte’s Wheel Theorem
Suppose for a contradiction that $G$ has a totally-nested nontrivial tri-separation $(A,B)$. Pick such an $(A,B)$ that is maximal with regard to the partial order $\leqslant$ on mixed-separations. Then $\sigma:=\{\,(A,B)\,\}$ is a splitting star of the set of totally-nested nontrivial tri-separations of $G$. Let $U:=B\setminus A$. By Section 2.7, the separator of $(A,B)$ consists of three vertices that do not span an edge. In particular, $N(U)=A\cap B$. Applying Section 2.7 together with Section 2.7 to $U$ yields that there is a strong nontrivial tri-separation $(C,D)$ of $G$ such that $S(C,D)$ contains a vertex of $U$ or an edge incident with a vertex of $U$. Since $(A,B)$ is totally-nested, this implies that $(A,B)<(C,D)$ or $(A,B)<(D,C)$. So $(C,D)$ interlaces $\sigma$. By Theorem 2.2.1, the compressed torso $X$ of $\sigma$ either is a wheel or a thickened $K_{3,m}$ or $G=K_{3,m}$ with $m\geqslant 0$. As $K_{3,m}$ has no totally-nested nontrivial tri-separation, $G$ is not a $K_{3,m}$. Note that as $S(A,B)$ consists only of vertices, $X$ is a genuine torso. Recall that $S(C,D)$ contains a vertex of $U$ or an edge.
@@
Automatic proof of Theorem 2.7.1.
Every edge of the graph $K_{3,m}$ with $m\geqslant 3$ is of type c. So by Section 2.7, $G$ is not a $K_{3,m}$ with $m\geqslant 3$. Applying Section 2.7 with $X=V(G)$ yields that $G$ has a nontrivial tri-separation, so $G$ is not internally 4-connected by Section 1.2. By Section 2.7, $G$ has no totally-nested nontrivial tri-separation. Hence by the (Angry Tri-Separation Theorem). (1.1.1), $G$ is a wheel. ∎
A natural problem in this area is to understand which edges of 3-connected graph are essential in that they cannot be contracted or deleted without destroying 3-connectivity; see for example [3], and [47] for further extensions. Theorem 2.2.1 and our automatic proof of Tutte’s wheel theorem provide a new perspective on essential edges, and it is not unreasonable to conjecture that these ideas can be used to resolve this problem.
Chapter 3 Concluding remarks
3.1. Tree-like decomposition
  1. (M1)
    $V(G)=\bigcup_{t\in T}V_{t}$;
  2. (M2)
We refer to the vertex sets $V_{t}$ as bags.
[…]
3.2. Outlook
@@
Extend Theorem 1.1.1 (and then Theorem 2.2.1) to 3-connected matroids.
@@
Extend Theorem 1.1.1 (and then Theorem 2.2.1) to separators of size larger than 3.
An instructive example concerning Section 3.2 is the line-graph of the 3-dimensional cube. In this graph, there are three 4-separations that cross ‘3-dimensionally’, as depicted in Figure 3.1 below.
Figure 3.1. Three 4-separations crossing 3-dimensionally
(Mixed) $k$-separations of order $k<4$ do not cross ‘3-dimensionally’: this is trivial for $k=1$; for $k=2$ we can read it from the (2-Separation Theorem). (1.6.1); and for $k=3$ this follows by combining the Nested or crossed: analysing corner diagrams (1.3) with Section 1.7. Some hope towards a solution of Section 3.2 stems from the results of [8] (building on earlier work of [64]), where it is proved that if a $k$-connected but not $(k+1)$-connected graph has minimum degree larger than $\frac{3k}{2}-1$, then it has a totally-nested $k$-separation (and in fact every $k$-separation $(A,B)$ with $A$ minimal is totally-nested).
@@
We expect that many results about 3-connected graphs in the literature can be derived in a fairly straightforward way from Theorem 2.2.1, for example those in [21] or [51].
Our main result Theorem 2.2.1 has quite a few applications in addition to the ones presented here. Whilst for some of these applications, our papers are at an early stage, the following applications will appear on the arXiv shortly (or already have appeared):
@@
In the following, we compare the decomposition of this paper with Grohe’s [38] and with the findings of the upcoming work [15, 14]. Details that we skip here will be addressed in [15, 14]. The results of this paper and related works give rise to three types of decompositions of 3-connected graphs, labelled below by (D1) to (D3). The decomposition (D1) is obtained by taking an inclusion-wise maximal set of pairwise nested nontrivial 3-separations; this is essentially the decomposition constructed by Grohe [38] and we refer to the upcoming work [15] for a refined analysis of this decomposition. The decomposition (D3) is that of Theorem 2.2.1. The decomposition (D2) is obtained from (D3) by applying to each quasi 4-connected compressed torso the decomposition (D1); so (D2) refines (D3).
@@
[…]
@@
Acknowledgements
We are grateful to Nathan Bowler and Reinhard Diestel for questions that indirectly led us to a streamlining of the definition of tri-separations. Raphael W. Jacobs and Paul Knappe have studied mixed-separations independently from us, and we thank them for interesting discussions on mixed-separations. We thank Dominik Blankenhagen and Nicolás Pich Preuss for spotting typos in the appendix.
@@ Acknowledgements
${}^{2}$ A set $X=\{v_{1},v_{2},v_{3},v_{4}\}$ of four vertices of a graph $G$ is cubic if there are components $C_{1},C_{2},C_{3},C_{4}$ of $G\setminus X$ such that $N(C_{i})=X-v_{i}$ for $i=1,2,3,4$, and no component of $G\setminus X$ has the whole of $X$ in its neighbourhood. If $X$ is cubic, then $G$ has the cube $Q_{3}$ as a minor where one bipartition class is $X$ and the other is $\{C_{1},C_{2},C_{3},C_{4}\}$. A 4-tangle is cubic if it lives on a cubic vertex set $X$; that is, every big side in the tangle contains $X$.
3.3. Appendix: Reviewing the 2-Separation Theorem
3.3.1. Overview of this chapter
A basic fact about graphs states that every connected graph can be cut along its cutvertices in a tree-like way into maximal 2-connected subgraphs and bridges. 2-connected graphs can be decomposed further in the same vein, which is useful to study planar embeddings of graphs, but it is no longer obvious where to best cut these graphs. MacLane found for every 2-connected graph $G$ a tree-decomposition of adhesion two all whose torsos are 3-connected, cycles or $K_{2}$’s [49]. Tutte [63] later found a canonical such tree-decomposition, for which Cunningham and Edmonds discovered an elegant one-step construction [20]. Here we review and prove a structural version of Tutte’s result with the description by Cunningham and Edmonds, using the terminology of this paper, and then derive the (2-Separation Theorem). (1.6.1) from it.
3.3.2. Characterising nestedness through connectivity
@@
The two situations in (X1) and (X2) of the following lemma are depicted in Figure 3.2. Recall that a separation $(A,B)$ separates two vertices $u,v$ if $u\in A\setminus B$ and $v\in B\setminus A$ or vice versa.
Figure 3.2. The two ways in which 2-separations can cross
@@
Case $X\cap Y=\emptyset$. We have to show that $(A,B)$ and $(C,D)$ cross like in a cycle. Let us suppose for a contradiction that they don’t. Then the two vertices in $X$, say, are not separated by $(C,D)$. With $X\cap Y=\emptyset$, it follows that $X\subseteq C\setminus D$, say. As $G[D]$ is connected by Section 3.3.2 and avoids $X$, it follows that $G[D]$ is included in a unique component $I$ of $G-X$. Without loss of generality, $I\subseteq G[B]$, so $B\supseteq D$. From $Y\subseteq I\subseteq G[B]$ and $X\cap Y=\emptyset$ we deduce that $G[A]$, which is connected by Section 3.3.2, is a connected subgraph of $G-Y$, and lies in the component $J$ of $G-Y$ that contains the subset $X\subseteq A$. Thus $G[A]\subseteq J\subseteq G[C]$, that is, $A\subseteq C$. Hence $(A,B)\leqslant(C,D)$, a contradiction.
Case $|X\cap Y|=1$. We will show that this case is impossible. Let us denote the vertex in the intersection $X\cap Y$ by $z$, and let us denote the vertices in $X\setminus Y$ and $Y\setminus X$ by $x$ and $y$, respectively; so $X=\{x,z\}$ and $Y=\{y,z\}$. Let $K(X,y)$ denote the component of $G-X$ that contains $y$, and let $K(Y,x)$ denote the component of $G-Y$ that contains $x$. Without loss of generality, $K(X,y)\subseteq G[B]$ and $K(Y,x)\subseteq G[C]$. Since $G[A]-z$ is connected by Section 3.3.2 and a subgraph of $G-Y$ that contains $x$, it must be included in the component $K(Y,x)$ of $G-Y$ that contains $x$. Hence $G[A]\subseteq K(Y,x)\cup N(K(Y,x))\subseteq G[C]$. A symmetric argument shows $G[D]\subseteq G[B]$. So $(A,B)\leqslant(C,D)$, a contradiction.
@@
Section 3.3.2 implies the following characterisation of total nestedness through external connectivity:
@@
3.3.3. When all 2-separations are crossed
Theorem 3.3.1 (Angry 2-Separation Theorem).
@@ 3.3.3. When all 2-separations are crossed
Proof of Theorem 3.3.1.
Let $(A,B)$ be a 2-separation of $G$. We may choose $(A,B)$ so that $G[A\setminus B]$ is connected. Let $(C,D)$ be a 2-separation of $G$ that crosses $(A,B)$. Since $G[A\setminus B]$ is connected, $(C,D)$ must cross $(A,B)$ like in a cycle. Let $T_{1}$ and $T_{2}$ be the block graphs of $G[A]$ and of $G[B]$, respectively, where we use the definition of block graphs as in [21, §3.1].
@@
Let us assume for a contradiction that $G[A]$, say, has a 2-connected block $X$. Let $Y$ denote the union of all blocks of $G[A]$ and of $G[B]$ except $X$. Then $\{V(X),V(Y)\}$ is a 2-separation of $G$. By Section 3.3.2, it it totally nested as $X$ is 2-connected and $Y\setminus X$ is connected, a contradiction. ∎
@@
[27, Theorem 3] Every vertex-transitive finite connected graph either is 3-connected, a cycle, a $K_{2}$ or a $K_{1}$.
@@
Let us suppose now that $G$ is not 3-connected, so $G$ has a 2-separation. If every 2-separation of $G$ is crossed by another one, then $G$ is a cycle of length $\geqslant 4$ by the (Angry 2-Separation Theorem). (3.3.1). Otherwise $G$ has a totally-nested 2-separation. Let $O$ denote its orbit under the action of the automorphism group of $G$, and pick $(A,B)\in O$ such that $A$ is minimal. Pick any vertex $v\in A\setminus B$. By vertex-transitivity, there is $(C,D)\in O$ such that $v\in C\cap D$. Since $O$ is nested, and since $v$ obstructs both of $(A,B)\leqslant(C,D)$ and $(A,B)\leqslant(D,C)$, we have $(C,D)\leqslant(A,B)$ or $(D,C)\leqslant(A,B)$. Hence $C\subseteq A$ or $D\subseteq A$. As $v$ lies in $C\cap D$ but not in $B$, the inclusion $C\subseteq A$ or $D\subseteq A$ must be proper, contradicting the choice of $(A,B)$. ∎
3.3.4. A structural 2-Separation Theorem
We say that $\sigma$ is $U$-principal for a vertex set $U\subseteq V(G)$ if $G\setminus U$ has at least three components and
@@ 3.3.4. A structural 2-Separation Theorem
Theorem 3.3.2 (Structural 2-Separation Theorem).
@@
We remark that the set $N$ in Theorem 3.3.2 is canonical.
@@
Clearly, $\sigma$ is a star. Its elements are totally nested by Section 3.3.2. Let $(C,D)$ be any 2-separation of $G$ that interlaces $\sigma$. Then $(C,D)$ defines a bipartition $(\mathscr{C},\mathscr{D})$ of the set of components of $G\setminus U$, where $\mathscr{C}$ and $\mathscr{D}$ consist of the components contained in $G[C]$ and in $G[D]$, respectively; also $C\cap D\subseteq U$, and hence $C\cap D=U$. Since $(C,D)$ is not in $\sigma$, both $\mathscr{C}$ and $\mathscr{D}$ contain at least two components. Hence $(C,D)$ is not totally-nested by Section 3.3.2. ∎
@@
Suppose for a contradiction that $(A_{0},B_{0})<(A_{1},B_{1})<\ldots<(A_{\omega},B_{\omega})$ is an $(\omega+1)$-chain in $N$. Since the elements of $N$ are half-connected, the separations in any 3-chain in $N$ do not all have the same separators. Therefore, we may assume without loss of generality that the $(A_{i},B_{i})$ have pairwise distinct separators. Let $x\in A_{0}\setminus B_{0}$ and $y\in B_{\omega}\setminus A_{\omega}$. Then $x$ and $y$ are separated by infinitely many pairwise distinct separators of size two. Since $G$ is 2-connected, these separators are inclusionwise minimal $x$$y$ separators in $G$. This contradicts a lemma of Halin [41, 2.4], which states that any two vertices $u,v$ in a graph are separated by only finitely many inclusionwise minimal $u$$v$ separators of size at most an arbitrarily prescribed $k\in\mathbb{N}$. The same argument also shows that $\bigcap_{n\in\mathbb{N}}\,(B_{n}\setminus A_{n})=\emptyset$ for every $\omega$-chain $(A_{0},B_{0})<(A_{1},B_{1})<\ldots$ in $N$. ∎
@@
form a star $\sigma\subseteq N$ that is interlaced by $(A,B)$. By Section 3.3.4, the star $\sigma$ is a splitting star of $N$. By Section 2.2.2, $(A,B)$ interlaces no other splitting star of $N$. ∎
Proof of Theorem 3.3.2.
@@
We claim that $X$ is 2-connected, and assume for a contradiction that it is not. Then $X$ has a separation $(A,B)$ of order at most one. By the Proof of (iii) (2.6), $(A,B)$ lifts to a separation $(\hat{A},\hat{B})$ of $G$ of order at most one, contradicting that $G$ is 2-connected. So $X$ is 2-connected.
@@
(i). Suppose that $\sigma$ is not interlaced by a 2-separation of $G$. If $X$ is not 3-connected, then $X$ has a 2-separation $(A,B)$ (since $X$ has at least four vertices). By the Proof of (iii) (2.6), $(A,B)$ lifts to a 2-separation $(\hat{A},\hat{B})$ of $G$, where it interlaces $\sigma$, a contradiction.
@@
Proof of Claim. Since $(A,B)$ interlaces $\sigma$, it is not in $N$. Hence $(A,B)$ is crossed by a 2-separation $(C,D)$ of $G$. We note that $(C,D)$ interlaces $\sigma$ as well, since otherwise $(C,D)$ would be nested with $(A,B)$. So the separators $A\cap B$ and $C\cap D$ are included in $X$. If $(A,B)$ and $(C,D)$ cross like in a cycle, then $(A\cap V(X),B\cap V(X))$ and $(C\cap V(X),D\cap V(X))$ are two crossing 2-separations of $X$ as desired. So it remains to show that $(A,B)$ and $(C,D)$ cannot cross with a four-flip. Indeed, otherwise $A\cap B=C\cap D$, and $G\setminus(A\cap B)$ has at least four components which define a splitting star of $N$ as in Section 3.3.4. As this splitting star is interlaced by $(A,B)$, it must be equal to $\sigma$ by Section 2.2.2. But then $V(X)=A\cap B$ contradicts our assumption that $X$ has at least four vertices. $\diamondsuit$
We recall that $X$ is 2-connected. By Claim 3.3.2.1 and our assumption, $X$ has a 2-separation. Every 2-separation $(A,B)$ of $X$ lifts to a 2-separation of $G$ by the Proof of (iii) (2.6), which interlaces $\sigma$ and through Claim 3.3.2.1 yields a 2-separation of $X$ that crosses $(A,B)$. Hence $X$ is a cycle of length $\geqslant 4$ by the (Angry 2-Separation Theorem). (3.3.1). ∎
3.3.5. Proof of the (2-Separation Theorem). (1.6.1)
@@
The proof of (i)$\to$(ii) is straightforward. The proof of [29, Lemma 2.7] shows (ii)$\to$(i), even though the statement of [29, Lemma 2.7] says otherwise. ∎
@@
Let $G$ be a 2-connected graph with a vertex $v$ of degree two such that $v$ lies in a 2-separator of $G$. Then, for every totally-nested 2-separation $(A,B)$ of $G$, we have $v\notin S(A,B)$ and $S(A,B)$ contains at most one neighbour of $v$.

### v3

@@ 7 Tutte’s Wheel Theorem
Suppose for a contradiction that $G$ has a totally-nested nontrivial tri-separation $(A,B)$. Pick such an $(A,B)$ that is maximal with regard to the partial order $\leqslant$ on mixed-separations. Then $\sigma:=\{\,(A,B)\,\}$ is a splitting star of the set of totally-nested nontrivial tri-separations of $G$. Let $U:=B\setminus A$. By Section 7, the separator of $(A,B)$ consists of three vertices that do not span an edge. In particular, $N(U)=A\cap B$. Applying Section 7 together with Section 7 to $U$ yields that there is a strong nontrivial tri-separation $(C,D)$ of $G$ such that $S(C,D)$ contains a vertex of $U$ or an edge incident with a vertex of $U$. Since $(A,B)$ is totally-nested, this implies that $(A,B)<(C,D)$ or $(A,B)<(D,C)$. So $(C,D)$ interlaces $\sigma$. By Theorem 2.1, the compressed torso $X$ of $\sigma$ either is a wheel or a thickened $K_{3,m}$ or $G=K_{3,m}$ with $m\geqslant 0$. As $K_{3,m}$ has no totally-nested nontrivial tri-separation, $G$ is not a $K_{3,m}$. Note that as $S(A,B)$ consists only of vertices, $X$ is a genuine torso. Recall that $S(C,D)$ contains a vertex of $U$ or an edge.
@@
Automatic proof of Theorem 7.1.
Every edge of the graph $K_{3,m}$ with $m\geqslant 3$ is of type c. So by Section 7, $G$ is not a $K_{3,m}$ with $m\geqslant 3$. Applying Section 7 with $X=V(G)$ yields that $G$ has a nontrivial tri-separation, so $G$ is not internally 4-connected by Section 3. By Section 7, $G$ has no totally-nested nontrivial tri-separation. Hence by the (Angry Tri-Separation Theorem). (1.1), $G$ is a wheel. ∎
A natural problem in this area is to understand which edges of 3-connected graph are essential in that they cannot be contracted or deleted without destroying 3-connectivity; see for example [3], and [49] for further extensions. Theorem 2.1 and our automatic proof of Tutte’s wheel theorem provide a new perspective on essential edges, and it is not unreasonable to conjecture that these ideas can be used to resolve this problem.
Chapter 3
Outlook
@@
Extend Theorem 1.1 (and then Theorem 2.1) to 3-connected matroids.
@@
Extend Theorem 1.1 (and then Theorem 2.1) to separators of size larger than 3.
Section 7 has recently been solved for size 4 in [51].
An instructive example concerning Section 7 is the line-graph of the 3-dimensional cube. In this graph, there are three 4-separations that cross ‘3-dimensionally’, as depicted in Figure 20 below.
Figure 20: Three 4-separations crossing 3-dimensionally
(Mixed) $k$-separations of order $k<4$ do not cross ‘3-dimensionally’: this is trivial for $k=1$; for $k=2$ we can read it from the (2-Separation Theorem). (7.1); and for $k=3$ this follows by combining the Nested or crossed: analysing corner diagrams (4) with Section 8. Some hope towards a solution of Section 7 stems from the results of [9] (building on earlier work of [67]), where it is proved that if a $k$-connected but not $(k+1)$-connected graph has minimum degree larger than $\frac{3k}{2}-1$, then it has a totally-nested $k$-separation (and in fact every $k$-separation $(A,B)$ with $A$ minimal is totally-nested).
@@
We expect that many results about 3-connected graphs in the literature can be derived in a fairly straightforward way from Theorem 2.1, for example those in [24] or [54].
Our main result Theorem 2.1 has quite a few applications in addition to the ones presented here. Whilst for some of these applications, our papers are at an early stage, the following applications will appear on the arXiv shortly (or already have appeared):
@@
In the following, we compare the decomposition of this paper with Grohe’s [39] and with the findings of the upcoming work [17, 15]. Details that we skip here will be addressed in [17, 15]. The results of this paper and related works give rise to three types of decompositions of 3-connected graphs, labelled below by (D1) to (D3). The decomposition (D1) is obtained by taking an inclusion-wise maximal set of pairwise nested nontrivial 3-separations; this is essentially the decomposition constructed by Grohe [39] and we refer to the upcoming work [17] for a refined analysis of this decomposition. The decomposition (D3) is that of Theorem 2.1. The decomposition (D2) is obtained from (D3) by applying to each quasi 4-connected compressed torso the decomposition (D1); so (D2) refines (D3).
@@
$\times$ $\times$ $\times$ (A1) 4-tangles appear1 as torsos
$\times$ (A2) non-cubic2 4-tangles appear as torsos
@@
2 A set $X=\{v_{1},v_{2},v_{3},v_{4}\}$ of four vertices of a graph $G$ is cubic if there are components $C_{1},C_{2},C_{3},C_{4}$ of $G\setminus X$ such that $N(C_{i})=X-v_{i}$ for $i=1,2,3,4$, and no component of $G\setminus X$ has the whole of $X$ in its neighbourhood. If $X$ is cubic, then $G$ has the cube $Q_{3}$ as a minor where one bipartition class is $X$ and the other is $\{C_{1},C_{2},C_{3},C_{4}\}$. A 4-tangle is cubic if it lives on a cubic vertex set $X$; that is, every big side in the tangle contains $X$.
Chapter 4
Reviewing the 2-Separation Theorem
1 Overview of this chapter
A basic fact about graphs states that every connected graph can be cut along its cutvertices in a tree-like way into maximal 2-connected subgraphs and bridges. 2-connected graphs can be decomposed further in the same vein, which is useful to study planar embeddings of graphs, but it is no longer obvious where to best cut these graphs. MacLane found for every 2-connected graph $G$ a tree-decomposition of adhesion two all whose torsos are 3-connected, cycles or $K_{2}$’s [52]. Tutte [66] later found a canonical such tree-decomposition, for which Cunningham and Edmonds discovered an elegant one-step construction [21]. Here we review and prove a structural version of Tutte’s result with the description by Cunningham and Edmonds, using the terminology of this paper, and then derive the (2-Separation Theorem). (7.1) from it.
2 Characterising nestedness through connectivity
@@
The two situations in (X1) and (X2) of the following lemma are depicted in Figure 21. Recall that a separation $(A,B)$ separates two vertices $u,v$ if $u\in A\setminus B$ and $v\in B\setminus A$ or vice versa.
Figure 21: The two ways in which 2-separations can cross
@@
Case $X\cap Y=\emptyset$. We have to show that $(A,B)$ and $(C,D)$ cross like in a cycle. Let us suppose for a contradiction that they don’t. Then the two vertices in $X$, say, are not separated by $(C,D)$. With $X\cap Y=\emptyset$, it follows that $X\subseteq C\setminus D$, say. As $G[D]$ is connected by Section 2 and avoids $X$, it follows that $G[D]$ is included in a unique component $I$ of $G-X$. Without loss of generality, $I\subseteq G[B]$, so $B\supseteq D$. From $Y\subseteq I\subseteq G[B]$ and $X\cap Y=\emptyset$ we deduce that $G[A]$, which is connected by Section 2, is a connected subgraph of $G-Y$, and lies in the component $J$ of $G-Y$ that contains the subset $X\subseteq A$. Thus $G[A]\subseteq J\subseteq G[C]$, that is, $A\subseteq C$. Hence $(A,B)\leqslant(C,D)$, a contradiction.
Case $|X\cap Y|=1$. We will show that this case is impossible. Let us denote the vertex in the intersection $X\cap Y$ by $z$, and let us denote the vertices in $X\setminus Y$ and $Y\setminus X$ by $x$ and $y$, respectively; so $X=\{x,z\}$ and $Y=\{y,z\}$. Let $K(X,y)$ denote the component of $G-X$ that contains $y$, and let $K(Y,x)$ denote the component of $G-Y$ that contains $x$. Without loss of generality, $K(X,y)\subseteq G[B]$ and $K(Y,x)\subseteq G[C]$. Since $G[A]-z$ is connected by Section 2 and a subgraph of $G-Y$ that contains $x$, it must be included in the component $K(Y,x)$ of $G-Y$ that contains $x$. Hence $G[A]\subseteq K(Y,x)\cup N(K(Y,x))\subseteq G[C]$. A symmetric argument shows $G[D]\subseteq G[B]$. So $(A,B)\leqslant(C,D)$, a contradiction.
@@
Section 2 implies the following characterisation of total nestedness through external connectivity:
@@
3 When all 2-separations are crossed
Theorem 3.1 (Angry 2-Separation Theorem).
@@ 3 When all 2-separations are crossed
Proof of Theorem 3.1.
Let $(A,B)$ be a 2-separation of $G$. We may choose $(A,B)$ so that $G[A\setminus B]$ is connected. Let $(C,D)$ be a 2-separation of $G$ that crosses $(A,B)$. Since $G[A\setminus B]$ is connected, $(C,D)$ must cross $(A,B)$ like in a cycle. Let $T_{1}$ and $T_{2}$ be the block graphs of $G[A]$ and of $G[B]$, respectively, where we use the definition of block graphs as in [24, §3.1].
@@
Let us assume for a contradiction that $G[A]$, say, has a 2-connected block $X$. Let $Y$ denote the union of all blocks of $G[A]$ and of $G[B]$ except $X$. Then $\{V(X),V(Y)\}$ is a 2-separation of $G$. By Section 2, it it totally nested as $X$ is 2-connected and $Y\setminus X$ is connected, a contradiction. ∎
@@
[28, Theorem 3] Every vertex-transitive finite connected graph either is 3-connected, a cycle, a $K_{2}$ or a $K_{1}$.
@@
Let us suppose now that $G$ is not 3-connected, so $G$ has a 2-separation. If every 2-separation of $G$ is crossed by another one, then $G$ is a cycle of length $\geqslant 4$ by the (Angry 2-Separation Theorem). (3.1). Otherwise $G$ has a totally-nested 2-separation. Let $O$ denote its orbit under the action of the automorphism group of $G$, and pick $(A,B)\in O$ such that $A$ is minimal. Pick any vertex $v\in A\setminus B$. By vertex-transitivity, there is $(C,D)\in O$ such that $v\in C\cap D$. Since $O$ is nested, and since $v$ obstructs both of $(A,B)\leqslant(C,D)$ and $(A,B)\leqslant(D,C)$, we have $(C,D)\leqslant(A,B)$ or $(D,C)\leqslant(A,B)$. Hence $C\subseteq A$ or $D\subseteq A$. As $v$ lies in $C\cap D$ but not in $B$, the inclusion $C\subseteq A$ or $D\subseteq A$ must be proper, contradicting the choice of $(A,B)$. ∎
4 A structural 2-Separation Theorem
We say that a star $\sigma$ is $U$-principal for a vertex set $U\subseteq V(G)$ if $G\setminus U$ has at least three components and
@@ 4 A structural 2-Separation Theorem
Theorem 4.1 (Structural 2-Separation Theorem).
@@
We remark that the set $N$ in Theorem 4.1 is canonical.
@@
Clearly, $\sigma$ is a star. Its elements are totally nested by Section 2. Let $(C,D)$ be any 2-separation of $G$ that interlaces $\sigma$. Then $(C,D)$ defines a bipartition $(\mathscr{C},\mathscr{D})$ of the set of components of $G\setminus U$, where $\mathscr{C}$ and $\mathscr{D}$ consist of the components contained in $G[C]$ and in $G[D]$, respectively; also $C\cap D\subseteq U$, and hence $C\cap D=U$. Since $(C,D)$ is not in $\sigma$, both $\mathscr{C}$ and $\mathscr{D}$ contain at least two components. Hence $(C,D)$ is not totally-nested by Section 2. ∎
@@
Suppose for a contradiction that $(A_{0},B_{0})<(A_{1},B_{1})<\ldots<(A_{\omega},B_{\omega})$ is an $(\omega+1)$-chain in $N$. Since the elements of $N$ are half-connected, the separations in any 3-chain in $N$ do not all have the same separators. Therefore, we may assume without loss of generality that the $(A_{i},B_{i})$ have pairwise distinct separators. Let $x\in A_{0}\setminus B_{0}$ and $y\in B_{\omega}\setminus A_{\omega}$. Then $x$ and $y$ are separated by infinitely many pairwise distinct separators of size two. Since $G$ is 2-connected, these separators are inclusionwise minimal $x$$y$ separators in $G$. This contradicts a lemma of Halin [42, 2.4], which states that any two vertices $u,v$ in a graph are separated by only finitely many inclusionwise minimal $u$$v$ separators of size at most an arbitrarily prescribed $k\in\mathbb{N}$. The same argument also shows that $\bigcap_{n\in\mathbb{N}}\,(B_{n}\setminus A_{n})=\emptyset$ for every $\omega$-chain $(A_{0},B_{0})<(A_{1},B_{1})<\ldots$ in $N$. ∎
@@
form a star $\sigma\subseteq N$ that is interlaced by $(A,B)$. To show that $\sigma$ is a splitting star of $N$, let $(X,Y)\in N$ be arbitrary. Without loss of generality we have $(X,Y)<(A,B)$. The set of all $(C,D)\in N$ with $(X,Y)\leqslant(C,D)<(A,B)$ has a maximal element by Section 4. This element is also contained in $\sigma$, so $\sigma$ is indeed a splitting star of $N$. By Section 2, $(A,B)$ interlaces no other splitting star of $N$. ∎
Proof of Theorem 4.1.
@@
We claim that $X$ is 2-connected, and assume for a contradiction that it is not. Then $X$ has a separation $(A,B)$ of order at most one. By the Proof of (iii) (6), $(A,B)$ lifts to a separation $(\hat{A},\hat{B})$ of $G$ of order at most one, contradicting that $G$ is 2-connected. So $X$ is 2-connected.
@@
(i). Suppose that $\sigma$ is not interlaced by a 2-separation of $G$. If $X$ is not 3-connected, then $X$ has a 2-separation $(A,B)$ (since $X$ has at least four vertices). By the Proof of (iii) (6), $(A,B)$ lifts to a 2-separation $(\hat{A},\hat{B})$ of $G$, which interlaces $\sigma$, a contradiction.
@@
Proof of Claim. Since $(A,B)$ interlaces $\sigma$, it is not in $N$. Hence $(A,B)$ is crossed by a 2-separation $(C,D)$ of $G$. We note that $(C,D)$ interlaces $\sigma$ as well, since otherwise $(C,D)$ would be nested with $(A,B)$. So the separators $A\cap B$ and $C\cap D$ are included in $X$. If $(A,B)$ and $(C,D)$ cross like in a cycle, then $(A\cap V(X),B\cap V(X))$ and $(C\cap V(X),D\cap V(X))$ are two crossing 2-separations of $X$ as desired. So it remains to show that $(A,B)$ and $(C,D)$ cannot cross with a four-flip. Indeed, otherwise $A\cap B=C\cap D$, and $G\setminus(A\cap B)$ has at least four components which define a splitting star of $N$ as in Section 4. As this splitting star is interlaced by $(A,B)$, it must be equal to $\sigma$ by Section 2. But then $V(X)=A\cap B$ contradicts our assumption that $X$ has at least four vertices. $\diamondsuit$
We recall that $X$ is 2-connected. By Claim 4.1.1 and our assumption, $X$ has a 2-separation. Every 2-separation $(A,B)$ of $X$ lifts to a 2-separation of $G$ by the Proof of (iii) (6), which interlaces $\sigma$ and through Claim 4.1.1 yields a 2-separation of $X$ that crosses $(A,B)$. Hence $X$ is a cycle of length $\geqslant 4$ by the (Angry 2-Separation Theorem). (3.1). ∎
5 Proof of the (2-Separation Theorem). (7.1)
@@
(1$\to$(2). Every $\omega$-chain in $N$ determines a ray in the decomposition tree $T$ of ${{\mathcal{T}}}(N)$. An $(\omega+1)$-chain in $N$ would determine a ray in $T$ followed by an edge, which is impossible. An $\omega$-chain in $N$ such as $(A_{0},B_{0})<(A_{1},B_{1})<\ldots$ determines a ray $R\subseteq T$, and a vertex $v\in\bigcap_{n\in\mathbb{N}}(B_{n}\setminus A_{n})$ would have to live in the direction of $R$ but in no bag in particular, contradicting (T1).
The proof of [30, Lemma 2.7] shows (2$\to$(1), even though the statement of [30, Lemma 2.7] says otherwise. ∎
@@
Let $G$ be a 2-connected graph with a vertex $v$ of degree two such that $v$ lies in a 2-separator of $G$. Then, for every totally-nested 2-separation $(A,B)$ of $G$, we have $v\notin S(A,B)$.